[Brahms-l] Shift report 20050611 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Jun 11 2005 - 08:03:25 EDT
Shifter leader was priyanka saxena
Shifters were 


1)Dump scheduled for 0100 and was injected successfully but was unexpectedly aborted because of RHIC VACUUM PROBLEM .<br>
2)Beam was restored at 5:13am.Since then we have it.Taken runs from 15383-15399.<br>
3)No detector problem as such.But everytime when TPCs&DCs were ON,A3_XU shows as tripped down.So I ahve to OFF it.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jun 11 08:04:10 2005

This archive was generated by hypermail 2.1.8 : Sat Jun 11 2005 - 08:04:17 EDT